Session: Oct 2009 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Catani at Mon Oct 26 20:05:44 2009) Success :

Yebes (Pablo de Vicente at Mon Oct 26 21:17:09 2009) Success :
No known problems

Onsala (Michael Lindqvist at Tue Oct 27 08:23:41 2009) Success :
Some rain. Otherwise no known problems.

Effelsberg (Uwe Bach at Tue Oct 27 08:26:24 2009) Minor failures :
Started 50 minutes too late, because of a human day-light saving time time lag!

Metsahovi (Guifre at Wed Oct 28 09:21:14 2009) Success :

Noto (P. Cassaro at Fri Oct 30 13:30:30 2009) Minor failures : RECO
Scan lost from 15:45 to 15:53 due to operator problems. The scans for the ftp fringe test have been extracted by hand (same problems as in the previous fringe test).

Jodrell_Bank (Alastair Gunn at Wed Nov 18 14:15:21 2009) Minor failures : RECO
The MkII telescope was used (Jb2). Several different disk packs tried at start of experiment, one of which failed, one which was throttling. No other known problems.

Cambridge (Alastair Gunn at Wed Nov 18 14:16:07 2009) Success :
No known problems.